during a competition a person applies a force of 2500 N to push a car 10 m. They are able to accomplish this in 1.3 minutes. How much power was exerted during this effort?


1
Expert's answer
2021-02-01T15:04:11-0500
"P=\dfrac{W}{t}=\dfrac{Fd}{t}=\dfrac{2500\ N\cdot10\ m}{1.3\ min\cdot\dfrac{60\ s}{1\ min}}=320.5\ W."